OpenStudy (bigbosssaint21):

What are the approximate solutions of 4x^2 + 2x = 17, rounded to the nearest hundredth? x ≈ -2.33 and x ≈ 1.83 x ≈ -1.83 and x ≈ 2.33 x ≈ -9.31 and x ≈ 7.31 No real solutions

OpenStudy (jdoe0001):

have you covered the quadratic formula yet?

OpenStudy (bigbosssaint21):

no lol

OpenStudy (jdoe0001):

hmm ok so... have you covered quadratic equations yet?

OpenStudy (bigbosssaint21):

some what but im not to good with it

OpenStudy (jdoe0001):

well... ok... the quadratic formula is quite simple \(\bf 4x^2 + 2x = 17\implies 4x^2 + 2x-17=0 \\ \quad \\ \textit{quadratic formula}\\ y={\color{blue}{ 4}}x^2{\color{red}{ +2}}x{\color{green}{ -17}} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (phi):

for more info, when you have time, see http://www.khanacademy.org/math/algebra/quadratics/quadratic-formula/v/using-the-quadratic-formula

OpenStudy (dan815):

so colorful..

OpenStudy (mathmate):

or use completing square: \[4x^2 + 2x = 17\] \[x^2 + x/2 = 17/4\] \[(x+1/4)^2 = 17/4+1/16\] so approximate solution is \[-\frac{1}{4}\pm 2\]